# Orchestrated Robust Controller for the Precision Control of Heavy-duty Hydraulic Manipulators

### v1

@@ V-B Hydraulic actuator dynamics and control
$f_{f}=Y_{f}\,\phi_{f}$ (86)
where $Y_{f}\in\mathbb{R}^{1\times 7}$ is the regressor matrix and $\phi_{f}\in\mathbb{R}^{7}$ is the friction parameter vector. The friction model in (86) includes coulomb friction, Stribeck friction, viscous friction, and the average deformation of the seal bristles, as well as a smooth transition between the presliding and sliding motions.
@@
$f_{p}=A_{a}\,p_{a}-A_{b}\,p_{b}$ (88)
with $A_{a}$ and $A_{b}$ being cross-sectional areas, and $p_{a}$ and $p_{b}$ being pressures in the hydraulic actuator chamber. The cylinder chamber pressure can be generated by controlling the fluid flow rate entering the chamber. The equations of these fluid flow rates, denoted as $Q_{a}$ and $Q_{b}$ can be written as:
$Q_{a}=c_{p1}\,\upsilon(p_{s}-p_{a})\,u\,\mathcal{S}(u)+c_{n1}\,\upsilon(p_{a}-p_{r})\,u\,\mathcal{S}(-u),$ (89)
$Q_{b}=-c_{n2}\,\upsilon(p_{b}-p_{r})\,u\,\mathcal{S}(u)-c_{p2}\,\upsilon(p_{s}-p_{b})\,u\,\mathcal{S}(-u),$ (90)
where $c_{p1}$, $c_{p2}$, $c_{n1}$ and $c_{n2}$ are flow coefficients, $p_{s}$ is the supply pressure, $p_{r}$ is the return-line pressure, and $\mathcal{S}(u)$ is the selection function, defined as:
$\mathcal{S}(u)=\left\{\begin{array}[]{ll}1,&\text{if}\,\quad u>0,\\ 0,&\text{if}\,\quad u\leqslant 0,\\ \end{array}\right.$ (93)
and $\upsilon(\Delta p)$ is the function related to drops in pressure:
$\upsilon(\Delta p)=\text{sign}(\Delta p)\,\sqrt{\left\lvert\Delta p\right\rvert}.$ (94)
@@
where $\beta$ denotes the oil bulk modulus, $s$ is the maximum stroke of the piston, and the laminar leakage flow $Q_{l}$ between the cylinder chambers is defined as:
$Q_{l}=c_{l}(p_{a}-p_{b}),$ (97)
with $c_{l}$ being the leakage coefficient. Taking the time derivative of the (88) and recalling (89), (90), and (95)-(97), one can obtain:
@@
where $Y_{v}=[Y_{v1},Y_{v2},Y_{v3},Y_{v4}]\in\mathbb{R}^{1\times 4}$ with $Y_{v1}=-\dfrac{\upsilon(p_{s}-p_{a})}{V_{0a}/A_{a}+x}\mathcal{S}(u)u$, $Y_{v2}=-\dfrac{\upsilon(p_{s}-p_{b})}{V_{0b}/A_{b}+(s-x))}\mathcal{S}(-u)u$, $Y_{v3}=-\dfrac{\upsilon(p_{a}-p_{r})}{V_{0a}/A_{a}+x}\mathcal{S}(-u)u$, and $Y_{v4}=-\dfrac{\upsilon(p_{b}-p_{r})}{V_{0b}/A_{b}+(s-x))}\mathcal{S}(u)u$, and $\theta_{v}=[c_{p1},c_{p2},c_{n1},c_{n2}]^{T}\in\mathbb{R}^{4}$. Additionally, $\Delta_{a}$ is the unknown model uncertainty in the actuator dynmaics.
@@
In order to ensure that the required piston force is generated at the actuator level, we need to design the voltage law $u_{fr}$. Considering (86), (87), (88), and (98) along with Lemma 1, we have:
@@
with $\theta_{d}=[1/\beta,A_{a},A_{b},c_{l}]^{T}\in\mathbb{R}^{4}$, and $Y_{d}=[Y_{d1},Y_{d2},Y_{d3},Y_{d4}]\in\mathbb{R}{{}^{1\times 4}}$ with $Y_{d1}=\dot{f}_{pr}$, $Y_{d2}=\dot{x}/(V_{0a}/A_{a}+x)$, $Y_{d3}=\dot{x}/(V_{0b}/A_{b}+(s-x))$, and $Y_{d4}=((pa-p_{b})(A_{a}V_{0b}+A_{b}V_{0a}+A_{a}A_{b}s))/((V_{0a}+A_{A}x)(V_{0b}+A_{b}(s-x)))$. Additionally, $\hat{W}_{a}\in\mathbb{R}^{\bar{n}_{a}\times 1}$ and $\hat{\varepsilon}_{a}\in\mathbb{R}$ are the estimation of $W_{a}$ and $\varepsilon_{a}=\varepsilon_{a}^{*}+\eta_{0}$ in the sense of Lemma 1 to handle the unmodeled dynamics in the hydraulic actuator. $\eta_{0}$ is defined in (85) and $\bar{n}_{a}$ is the total number of neurons in the actuator subsystem. The signal in (102) is the required voltage that accomplishes the control goals in the absence of deadzone-backlash, while $v_{fr}$ is equipped with the input constraint compensator in the sense of (80) as:
@@
$u_{fr}=-Y_{v}\hat{\theta}_{v}$ (106)
with $\hat{(.)}$ being the estimation of (.)(.)( . ). The low-level control law designed in (103) not only compensates for compound input constraints and unknown unmodeled dynamics at the actuator level, but it also makes the actuator produce the required piston forces that tackle unknown model uncertainty and unknown disturbances in rigid body part, resulting in precise trajectory tracking.
@@ VI Stability Analysis
Consider the 6-DoF hydraulic manipulator demonstrated in Fig. 1 that decomposed into rigid body and actuator subsystems in Fig. 2-4. The rigid body dynamics (31) in the presence of unknown model uncertainty and disturbance under the robust local controller (67) with the adaptation law of:
@@
is virtually stable, where $A\in S$ and $S=\left\{P_{1},P_{p2},B_{0j},B_{1j},B_{3j},B_{4j},G_{i}\right\}$ for $i=1,2,3$ and $j=1,2$. Moreover, ^${}^{A}\hat{\mathcal{L}}$ is the estimation of ${}^{A}\mathcal{L}$ defined in Lemma 2 and ${}^{A}\mathcal{S}$ is a unique symmetric matrix. The $\gamma,\gamma_{0},\tau_{0},\pi,\pi_{0}$ are positive constants, and $\Gamma$ is a positive-definite matrix. Let the non-negative accompanying function for the rigid body part be chosen in the sense of Definition 3 as ${\nu}_{1}$. Then, the time derivative of ${\nu}_{1}$ is:
@@
is virtually stable, with (.)(.)( . ) denoting all the objects. Defining the non-negative accompanying functions for linear hydraulic actuators as $\nu_{a}$ in the sense of Definition 3 yields to:
@@
with $\mu_{a}$ and $\mu_{a0}$ being positive, and VPFs being the virtual cutting points at the actuator level.
@@
Consider the entire 6-DoF heavy-duty hydraulic manipulator with rigid body and actuator dynamics represented in (31)-(47) and (86)-(100) with control laws of (67), (74), (75), (76), (102), and (103) and adaptation functions in (107)-(109) and (111)-(116). All the signals and errors in the system are SGUUB and will remain within a compact set. For example, velocity tracking error, actuator force tracking error, and RBFNNs weight estimation error will stay within a compact set $\Omega_{V}$, $\Omega_{f}$, and $\Omega_{W}$, respectively, where $\Omega_{V}=\{e_{V}\in\mathbb{R}^{6}|\,||e_{v}||\leq\sqrt{\frac{2B}{\lambda_{min}(M)}}\}$, $\Omega_{f}=\{e_{f}\in\mathbb{R}|\,||e_{f}||\leq\sqrt{2B\beta k_{x}}\}$, and $\Omega_{W}=\{\tilde{W}_{a}\in\mathbb{R}^{n}|\,||\tilde{W}_{a}||\leq\sqrt{2B\delta_{x}}\}$ with $B=\bar{\mu}_{0}/\bar{\mu}$. Similarly, for all the errors, a compact set can be achieved.
@@ VII Results and Discussion
In this section, the simulation and experimental results are provided to evaluate the performance of the proposed controller. First, the simulation results of comparing the presented method with three different controllers, including the original VDC, proportional-differential (PD) controller, and the state-of-the-art controller, adaptive dynamic surface controller with funnel control and neural networks (ADSC-FC-NN) proposed in [16]. The simulation analysis is performed for the base joint of the HHM, shown in Fig. 1, which is subjected to deadzone in valves and backlash in the rack and pinion mechanism. Then, the performance of the proposed method is implemented to control the real-world commercial heavy-duty hydraulic manipulator. The RVDC in the figures is an abbreviation for robust VDC, which represents the presented method.
@@ VII-A Performance comparison in Simulation
In the simulation, which is performed by MATLAB & SIMULINK, the deadzone and backlash constraints are considered to evaluate the performance of each controller in the presence of such compound constraints, as well as model uncertainties. The deadzone and backlash parameters are set to $b_{r}=-b_{l}=0.2$, $m_{d}=1$, $B_{r}=-B_{l}=0.2$, $k_{b}=1$. All the control gains are tuned to achieve the best results for each controller, and the 10 nodes in total are utilized for RBFNNs in the proposed method and ADSC-FC-NN. Fig. 7 demonstrates the result of the simulation. The desired trajectory in Fig. 7a helps to analyze both the transient and steady-state performances of each controller. As displayed in the figure, the ADSC-FC-NN controller has much better transient performance than the original VDC and PD. However, both the VDC and ADSC-FC-NN could not converge to the desired point, and both oscillated around the set point, with errors of 0.17 and 0.25 degrees, respectively, due to the compound input constraint. This shows that the RBFNN in ADSC-FC-NN could not perfectly estimate the uncertainty in the actuator. In contrast, the proposed method not only improved the weakness of the original VDC in the transient part, but it also ultimately converged to the desired point and perfectly tackled the effect of input constraints. Fig. 7b compares the errors of each controller. It can be concluded that using novel DRBFNNs for the estimation of actuator uncertainties, which estimation of the non-parametrizable term in (104) is part of it, based on the states of the actuator, perfectly accomplished the control objective. Table I additionally provides a quantitative performance evaluation of each controller by comparing the maximum-absolute of tracking error $|e_{max}|$, root-mean-square (RMS) error (RMSE) $e_{rms}$, and RMS of computed voltage $u_{rms}$. According to Table I, the maximum error value of the ADSC-FC-NN is better than the proposed method, while the designed controller has a much lower RMS error and almost zero steady-state error with lower voltage usage.
@@
Controller
PD
ADSC-NN-FC [16]
VDC
RVDC
$|e|_{max}(^{\circ})$
4.57
$0.53$
1.88
0.6
$e_{rms}(^{\circ})$
1.04
$0.22$
0.46
0.12
$u_{rms}(v)$
0.15
$0.74$
0.34
0.44
@@ VII-B Experimental Set-up
Symbol
Value
Symbol
Value
$\lambda$
3
$\bar{\lambda}_{xj}$
5
$\sigma_{1,2}$
12
$\sigma_{3}$
18
$k_{xp}$
0.02
$k_{xj}$
0.015
$k_{xwi}$
0.08
$k_{fp}$
$0.01\cdot 10^{-8}$
$k_{fj}$
$2\cdot 10^{-8}$
$k_{fwi}$
$1\cdot 10^{-8}$
$K_{A}$
$50\cdot\boldsymbol{I}$
$\gamma$
$500$
${{}^{A}}\Gamma$
$350$
${{}^{A}}\pi$
$20$
$\delta_{a(.)}$
$1.5\cdot k_{x(.)}k_{f(.)}$
$\bar{\delta}_{a(.)}$
$0.5\cdot k_{x(.)}k_{f(.)}$
@@ VII-C Experimental Result
In this section, the experimental results of implementing the proposed controller on an industrial 6-DoF HHM are provided. The result of the proposed method is compared to that of the original VDC controller to evaluate the control performance more accurately. Table II provides the control gains utilized in experiments. For the RBFNNs, the Gaussian activation function is defined as $\Psi(\chi)=exp([-(\chi-c_{\mathbf{j}})^{T}(\chi-c_{\mathbf{j}})/(b_{\mathbf{j}}^{2})])$, with $c_{\mathbf{j}}$ and $b_{\mathbf{j}}$ denoting the center and width of the neural cell in $\mathbf{j}$th unit, with $\mathbf{j}=\bar{n}_{A}$ for rigid body subsystem, and $\mathbf{j}=\bar{n}_{a}$ for actuator subsystem. In this study the value for $c_{\mathbf{j}}$ is randomly selected in $[-1,1]$ with $b_{\bar{n}_{a}}=0.5$ and $b_{\bar{n}_{A}}=5$ with the aggregation of used 270 nodes.
@@
In Fig. 10, the performance of the proposed method and the original VDC in tracking a desired trajectory is compared. As demonstrated, both controllers performed excellently in following the desired trajectory. However, from Fig. 11, it can be concluded that the proposed controller achieved a much better result by handling the uncertainties in the model and considering the input constraints. As mentioned previously, the base and wrist actuators (objects 1 and 3) suffer from compound input constraints. Such uncertainty along with the model uncertainties lead to the steady-state error of 0.13, -0.6, -0.7, and 0.07 degree in $\zeta_{1}$, $\xi_{1}$, $\xi_{2}$, and $\xi_{3}$, respectively, using VDC controller. By employing the presented controller, the errors are reduced to -0.04, -0.01, -0.02, and -0.01 degrees, respectively. In addition to the steady state errors, the proposed method achieved a considerably lower transient error, showing the perfect performance of the designed method in tackling unknown uncertainties and compound input constraints. Fig. 12, on the other hand, depicts the angular velocity of each joint along with the desired values, displaying that the RVDC ended up with much smoother tracking without jumps in velocities.
@@
Fig. 10: Trajectory tracking of all six joints. The blue solid line is the desired trajectory, the dashed red line is with RVDC, and the black solid line is with VDC. The y-axis is in degrees.
@@
Fig. 11: Trajectory tracking error of each joint. The dashed red line is with RVDC, and the black solid line is with VDC. The y-axis is in degrees.
@@
Fig. 12: Angular velocities of all joints. The blue solid line is the desired trajectory, the dashed red line is with RVDC, and the black solid line is with VDC. The y-axis is in degrees/sec.
@@
Controllers
B
L
T
W1
W2
W3
$e_{rms}(^{\circ})$
VDC
0.14
0.15
0.12
0.94
1.38
0.59
RVDC
0.07
0.03
0.06
0.18
0.19
0.25
$|e|_{max}(^{\circ})$
VDC
0.4
0.62
0.38
1.77
2.6
1.3
RVDC
0.23
0.16
0.24
0.89
0.8
1.03
$\tilde{f}_{rms}(kN)$
VDC
15.3
2
1.13
31.53
44.82
31.5
RVDC
13.3
3.1
1.2
16.74
19.4
18.9
$u_{rms}(v)$
VDC
0.7
0.6
0.55
0.53
0.4
0.62
RVDC
0.51
0.56
0.55
0.51
0.41
0.62
The lower force error $(\tilde{f}=f_{cr}-f_{c})$, shows the better performance of the hydraulic actuator in generating the required forces. Such a good performance of the actuator can be dependent on friction and model uncertainties. The experimental results of the actuator force tracking error are demonstrated in Fig. 13, from which can be seen that the actuator with RVDC had a much lower force error than VDC, displaying the impact of contributions of this study. Fig. 14, also, shows the time history of the estimated states. Fig. 14a shows the norm of all the estimated inertial parameter vectors of the rigid body in (67). Fig. 14b displays the estimated adaptive inverse deadzone-backlash parameters. Fig. 14c and 14d demonstrate the norm of estimation of weights and approximation errors of the DRBFNNs in rigid body subsystems. Finally, Fig. 14e and 14f, show the norm of estimated weights and approximation error in the actuator subsystems. All the presented figures show the boundedness of all signals and the good performance of the proposed method, validating the theoretical results. Table III summarizes the comparison between the original VDC and the proposed method. It can be seen that for a given roughly the same power source, the proposed approach achieved considerably lower tracking errors in comparison to the original VDC [12, 36]. The generated actuator forces, also, with RVDC have a lower difference with the required value in contrast to the original VDC.
Fig. 13: force errors of each actuator with the implemented controllers. The dashed red line is with RVDC, and the black solid line is with VDC. The y-axis is in kN.
Fig. 14: Time history of estimated parameters. a) norm of estimated inertial parameters, b) estimation of deadzone-backlash parameters, c) time history of updated weights of RBFNNs for the rigid body, d) estimation of bias error, e) weight estimation of RBFNNs for actuators, f) bias error estimation of actuators.
Fig. 15: Position and orientation tracking error of the end-effector in the $x-y-z$ plane. The solid black line is the tracking error with VDC and the dashed red line is with RVDC. The position and orientation errors are presented in $cm$ and $deg$, respectively.

### v2

@@ V-B Hydraulic actuator dynamics and control
$f_{f}=Y_{f}\,\phi_{f}.$ (84)
The friction model in (84) includes coulomb friction, Stribeck friction, viscous friction, and the average deformation of the seal bristles, as well as a smooth transition between the presliding and sliding motions.
@@
$f_{p}=A_{a}\,p_{a}-A_{b}\,p_{b}.$ (86)
The cylinder chamber pressure can be generated by controlling the fluid flow rate entering the chamber. The equations of these fluid flow rates, denoted as $Q_{a}$ and $Q_{b}$ can be written as:
$Q_{a}=c_{p1}\,\mathcal{Z}(p_{s}-p_{a})\,u\,\mathcal{W}(u)+c_{n1}\,\mathcal{Z}(p_{a}-p_{r})\,u\,\mathcal{W}(-u),$ (87)
$Q_{b}=-c_{n2}\,\mathcal{Z}(p_{b}-p_{r})\,u\,\mathcal{W}(u)-c_{p2}\,\mathcal{Z}(p_{s}-p_{b})\,u\,\mathcal{W}(-u),$ (88)
where $\mathcal{W}(u)$ is the selection function, defined as:
$\mathcal{W}(u)=\left\{\begin{array}[]{ll}1,&\text{if}\,\quad u>0,\\ 0,&\text{if}\,\quad u\leqslant 0,\\ \end{array}\right.$ (91)
and $\mathcal{Z}(\Delta p)$ is the function related to drops in pressure:
$\mathcal{Z}(\Delta p)=\text{sign}(\Delta p)\,\sqrt{\left\lvert\Delta p\right\rvert}.$ (92)
@@
where the laminar leakage flow $Q_{l}$ between the cylinder chambers is defined as:
$Q_{l}=c_{l}(p_{a}-p_{b}).$ (95)
Taking the time derivative of the (86) and recalling (87), (88), and (93)-(95), one can obtain:
@@
where $Y_{v}=[Y_{v1},Y_{v2},Y_{v3},Y_{v4}]$ with $Y_{v1}=-\dfrac{\upsilon(p_{s}-p_{a})}{V_{0a}/A_{a}+x}\mathcal{S}(u)u$, $Y_{v2}=-\dfrac{\upsilon(p_{s}-p_{b})}{V_{0b}/A_{b}+(s-x))}\mathcal{S}(-u)u$, $Y_{v3}=-\dfrac{\upsilon(p_{a}-p_{r})}{V_{0a}/A_{a}+x}\mathcal{S}(-u)u$, and $Y_{v4}=-\dfrac{\upsilon(p_{b}-p_{r})}{V_{0b}/A_{b}+(s-x))}\mathcal{S}(u)u$, and $\theta_{v}=[c_{p1},c_{p2},c_{n1},c_{n2}]^{T}$.
@@
In order to ensure that the required piston force is generated at the actuator level, we need to design the voltage law $u_{fr}$. Considering (84), (85), (86), and (96) along with Definition 6, we have:
@@
with $\theta_{d}=[1/\beta,1,1,c_{l}]^{T}$, and $Y_{d}=[Y_{d1},Y_{d2},Y_{d3},Y_{d4}]$ with $Y_{d1}=\dot{f}_{pr}$, $Y_{d2}=(A_{a}\dot{x})/(V_{0a}/A_{a}+x)$, $Y_{d3}=(A_{b}\dot{x})/(V_{0b}/A_{b}+(s-x))$, and $Y_{d4}=((pa-p_{b})(A_{a}V_{0b}+A_{b}V_{0a}+A_{a}A_{b}s))/((V_{0a}+A_{A}x)(V_{0b}+A_{b}(s-x)))$. Additionally, $\chi_{a}=[x,\dot{x},p_{a},p_{b}]^{T}\in\mathbb{R}^{4}$, $\hat{W}_{a}\in\mathbb{R}^{\bar{n}_{a}\times 1}$ and $\hat{\varepsilon}_{a}\in\mathbb{R}$ are the estimation of $W_{a}$ and $\varepsilon_{a}=\varepsilon_{a}^{*}+\eta_{0}$ in the sense of Definition 6 to handle the unmodeled dynamics in the hydraulic actuator. $\eta_{0}$ is defined in (83) and $\bar{n}_{a}$ is the total number of neurons in the actuator subsystem. The signal in (100) is the required voltage that accomplishes the control goals in the absence of deadzone-backlash, while $v_{fr}$ is equipped with the input nonlinearities compensator in the sense of (78) as:
@@
$u_{fr}=-Y_{v}\hat{\theta}_{v}.$ (104)
The low-level control law designed in (101) not only compensates for compound input nonlinearities and unknown unmodeled dynamics at the actuator level, but it also makes the actuator produce the required piston forces that tackle unknown model uncertainty and unknown disturbances in rigid body part, resulting in precise trajectory tracking.
@@ VI Stability Analysis
Consider the 6-DoF hydraulic manipulator demonstrated in Fig. 1 that decomposed into rigid body and actuator subsystems in Fig. 2-4. The rigid body dynamics (31) in the presence of unknown model uncertainty and disturbance under the robust local controller (65) with the adaptation laws of:
@@
is virtually stable in the sense of Definition 4, where $A\in S$ and $S=\left\{P_{1},P_{p2},B_{0j},B_{1j},B_{3j},B_{4j},G_{i}\right\}$ for $i=1,2,3$ and $j=1,2$. Moreover, ${}^{A}\mathcal{S}$ is a unique symmetric matrix. Let the non-negative accompanying function for the rigid body part be chosen in the sense of Definition 3 as ${\nu}_{1}$. Then, the time derivative of ${\nu}_{1}$ is:
@@
is virtually stable in the sense of Definition 4, with (.)(.)( . ) denoting all the objects. Defining the non-negative accompanying functions for linear hydraulic actuators as $\nu_{a}$ in the sense of Definition 3 yields to:
@@
with $\mu_{a}$ and $\mu_{a0}$ being positive, and VPFs being the virtual power flow at the actuator level.
@@
Consider the generic 6-DoF HHM with rigid body and actuator dynamics represented in (31)-(45) and (84)-(98) with control laws of (65), (72), (73), (74), (100), and (101) and adaptation functions in (105)-(107) and (109)-(114). All the signals and errors in the system are SGUUB and will remain within a compact set. For example, velocity tracking error, actuator force tracking error, and RBFNNs weight estimation error will stay within a compact set $\Omega_{V}$, $\Omega_{f}$, and $\Omega_{W}$, respectively, where $\Omega_{V}=\{e_{V}\in\mathbb{R}^{6}|\,||e_{v}||\leq\sqrt{\frac{2B}{\lambda_{min}(M)}}\}$, $\Omega_{f}=\{e_{f}\in\mathbb{R}|\,||e_{f}||\leq\sqrt{2B\beta k_{x}}\}$, and $\Omega_{W}=\{\tilde{W}_{a}\in\mathbb{R}^{n}|\,||\tilde{W}_{a}||\leq\sqrt{2B\delta_{x}}\}$ with $B=\bar{\mu}_{0}/\bar{\mu}$. Similarly, for all the errors, a compact set can be achieved.
@@
Remark 4
The adaptation laws in (105)-(107) and (109)-(114) are structured to achieve SGUUB. This ensures not only that the system remains bounded and robust but also that the control terms converge reliably under various uncertainties.
@@ VII Results and Discussion
In this section, the simulation and experimental results are provided to evaluate the performance of the proposed controller. First, the simulation results of comparing the presented ORC method with three different controllers, including the original VDC, proportional-differential (PD) controller, and the state-of-the-art controller, adaptive dynamic surface controller with funnel control and neural networks (ADSC-FC-NN) proposed in [17]. The simulation analysis is performed for the base joint of the HHM, shown in Fig. 1, which is subjected to deadzone in valves and backlash in the rack and pinion mechanism. Then, the performance of the proposed method is implemented to control the real-world commercial heavy-duty hydraulic manipulator.
@@ VII-A Performance comparison in Simulation
In the simulation, which is performed by MATLAB & SIMULINK, the deadzone and backlash constraints are considered to evaluate the performance of each controller in the presence of such compound constraints, as well as model uncertainties. The deadzone and backlash parameters are set to $b_{r}=-b_{l}=0.2$, $m_{d}=1$, $B_{r}=-B_{l}=0.2$, $k_{b}=1$. All the control gains are tuned to achieve the best results for each controller, and the 10 nodes in total are utilized for RBFNNs in the proposed method and ADSC-FC-NN. Fig. 7 demonstrates the result of the simulation. The desired trajectory in Fig. 7a helps to analyze both the transient and steady-state performances of each controller. As displayed in the figure, the ADSC-FC-NN controller has much better transient performance than the original VDC and PD. However, both the VDC and ADSC-FC-NN could not converge to the desired point, and both oscillated around the set point, with errors of 0.17 and 0.25 degrees, respectively, due to the compound input nonlinearities. This shows that the RBFNN in ADSC-FC-NN could not perfectly estimate the uncertainty in the actuator. In contrast, the proposed method not only improved the weakness of the original VDC in the transient part, but it also ultimately converged to the desired point and perfectly tackled the effect of input nonlinearities. Fig. 7b compares the errors of each controller. It can be concluded that using novel DRBFNNs for the estimation of actuator uncertainties, which estimation of the non-parametrizable term in (102) is part of it, based on the states of the actuator, perfectly accomplished the control objective. Table I additionally provides a quantitative performance evaluation of each controller by comparing the maximum-absolute of tracking error $|e_{max}|$, root-mean-square (RMS) error (RMSE) $e_{rms}$, and RMS of computed voltage $u_{rms}$. According to Table I, the maximum error value of the ADSC-FC-NN is better than the proposed method, while the designed controller has a much lower RMS error and almost zero steady-state error with lower voltage usage.
@@
Controller PD ADSC-NN-FC [17] VDC ORC
$|e|_{max}(^{\circ})$ 4.57 $0.53$ 1.88 0.6
$e_{rms}(^{\circ})$ 1.04 $0.22$ 0.46 0.12
$u_{rms}(v)$ 0.15 $0.74$ 0.34 0.44
@@ VII-B Experimental Set-up
Symbol Value Symbol Value
$\lambda$ 3 $\bar{\lambda}_{xj}$ 5
$\sigma_{1,2}$ 12 $\sigma_{3}$ 18
$k_{xp}$ 0.02 $k_{xj}$ 0.015
$k_{xwi}$ 0.08 $k_{fp}$ $0.01\cdot 10^{-8}$
$k_{fj}$ $2\cdot 10^{-8}$ $k_{fwi}$ $1\cdot 10^{-8}$
$K_{A}$ $50\cdot\boldsymbol{I}$ $\gamma$ $500$
${{}^{A}}\Gamma$ $350$ ${{}^{A}}\pi$ $20$
$\delta_{a(.)}$ $1.5\cdot k_{x(.)}k_{f(.)}$ $\bar{\delta}_{a(.)}$ $0.5\cdot k_{x(.)}k_{f(.)}$
@@ VII-C Experimental Result
In this section, the experimental results of implementing the proposed controller on an industrial 6-DoF HHM are provided. The result of the proposed method is compared to that of the original VDC controller to evaluate the control performance more accurately. Table II provides the control gains utilized in experiments. For the RBFNNs, the Gaussian activation function is defined as $\Psi(\chi)=exp([-(\chi-c_{\mathbf{j}})^{T}(\chi-c_{\mathbf{j}})/(b_{\mathbf{j}}^{2})])$, with $c_{\mathbf{j}}$ and $b_{\mathbf{j}}$ denoting the center and width of the neural cell in $\mathbf{j}$th unit, with $\mathbf{j}=\bar{n}_{A}$ for rigid body subsystem, and $\mathbf{j}=\bar{n}_{a}$ for actuator subsystem. In this study the value for $c_{\mathbf{j}}$ is randomly selected in $[-1,1]\times[-1,1]\times[-1,1]$ and $[-1,1]\times[-1,1]\times[-1,1]\times[-1,1]$ for $\chi_{A}$ and $\chi_{a}$, respectively, with $b_{\bar{n}_{a}}=0.5$ and $b_{\bar{n}_{A}}=5$ with the aggregation of 270 nodes. Additionally, prior to inputting the vectors into RBFNN, we normalize all input features to the range of $[-1,1]$. This step ensures that all input dimensions are treated uniformly and that the RBFNNs can effectively cover the input space without the risk of incompatibility.
@@
In Fig. 10, the performance of the proposed method and the original VDC in tracking a desired trajectory is compared. As demonstrated, both controllers performed excellently in following the desired trajectory. However, from Fig. 11, it can be concluded that the proposed controller achieved a much better result by handling the uncertainties in the model and considering the input nonlinearities. As mentioned previously, the base and wrist actuators (objects 1 and 3) suffer from compound input nonlinearities. Such uncertainty along with the model uncertainties lead to the steady-state error of 0.13, -0.6, -0.7, and 0.07 degree in $\zeta_{1}$, $\xi_{1}$, $\xi_{2}$, and $\xi_{3}$, respectively, using VDC controller. By employing the presented controller, the errors are reduced to -0.04, -0.01, -0.02, and -0.01 degrees, respectively. In addition to the steady state errors, the proposed method achieved a considerably lower transient error, showing the perfect performance of the designed method in tackling unknown uncertainties and compound input nonlinearities. Fig. 12, on the other hand, depicts the angular velocity of each joint along with the desired values, displaying that the ORC ended up with much smoother tracking without jumps in velocities.
@@
Fig. 10: Trajectory tracking of all six joints. The blue solid line is the desired trajectory, the dashed red line is with ORC, and the black solid line is with VDC. The y-axis is in degrees.
@@
Fig. 11: Trajectory tracking error of each joint. The dashed red line is with ORC, and the black solid line is with VDC. The y-axis is in degrees.
@@
Fig. 12: Angular velocities of all joints. The blue solid line is the desired trajectory, the dashed red line is with ORC, and the black solid line is with VDC. The y-axis is in degrees/sec.
The lower force error $(\tilde{f}=f_{cr}-f_{c})$, shows the better performance of the hydraulic actuator in generating the required forces. Such a good performance of the actuator can be dependent on friction and model uncertainties. The experimental results of the actuator force tracking error are demonstrated in Fig. 13, from which can be seen that the actuator with ORC had a much lower force error than VDC, displaying the impact of contributions of this study. Fig. 14, also, shows the time history of the estimated states. Fig. 14a shows the norm of all the estimated inertial parameter vectors of the rigid body in (65). Fig. 14b displays the estimated adaptive inverse deadzone-backlash parameters. Fig. 14c and 14d demonstrate the norm of estimation of weights and approximation errors of the DRBFNNs in rigid body subsystems. Finally, Fig. 14e and 14f, show the norm of estimated weights and approximation error in the actuator subsystems. All the presented figures show the boundedness of all signals and the good performance of the proposed method, validating the theoretical results. Table III summarizes the comparison between the original VDC and the proposed method. It can be seen that for a given roughly the same power source, the proposed approach achieved considerably lower tracking errors in comparison to the original VDC [12, 36]. The generated actuator forces, also, with ORC have a lower difference with the required value in contrast to the original VDC.
Fig. 13: force errors of each actuator with the implemented controllers. The dashed red line is with ORC, and the black solid line is with VDC. The y-axis is in kN.
Fig. 14: Time history of estimated parameters, in which, each line displays the components of the corresponding vectors. a) norm of estimated inertial parameters, b) estimation of deadzone-backlash parameters, c) time history of updated weights of RBFNNs for the rigid body, d) estimation of bias error, e) weight estimation of RBFNNs for actuators, f) bias error estimation of actuators.
@@
Controllers B L T W1 W2 W3
$e_{rms}(^{\circ})$ VDC 0.14 0.15 0.12 0.94 1.38 0.59
ORC 0.07 0.03 0.06 0.18 0.19 0.25
$|e|_{max}(^{\circ})$ VDC 0.4 0.62 0.38 1.77 2.6 1.3
ORC 0.23 0.16 0.24 0.89 0.8 1.03
$\tilde{f}_{rms}(kN)$ VDC 15.3 2 1.13 31.53 44.82 31.5
ORC 13.3 3.1 1.2 16.74 19.4 18.9
$u_{rms}(v)$ VDC 0.7 0.6 0.55 0.53 0.4 0.62
ORC 0.51 0.56 0.55 0.51 0.41 0.62
Fig. 15: Position and orientation tracking error of the end-effector in the $x-y-z$ plane. The solid black line is the tracking error with VDC and the dashed red line is with ORC. The position and orientation errors are presented in $cm$ and $deg$, respectively.